OpenStudy (anonymous):

To be considered safe for ocean sailing, the capsize screening value C should be less than 2 For a boat with a beam (or width) b in feet and displacement d in pounds, C is determined by the function C=4d^-1/3b. a) Find the capsize screening value for the Tartan 4100 which has a displacement of 23,245 lbs and a beam of 13.5 ft. ( I know how to set up the problem but not how to solve it I have an example of how to solve but I do not understand how they got the first part of the answer .. If I find out how I can solve it )) ANY help is greatly appreciated ! :)

OpenStudy (anonymous):

so.. you have 3 variables C, b, d. the question gives you b = 13.5 and d = 23,245, simply plug them in the equation to solve for C

OpenStudy (anonymous):

yeah i know how to set it all up but Im not sure how they got the answer they did for the first part in my example if I show you do you think you could help explain how they got the answer so I can solve mine ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[C=4d ^{-1/3}b\]

OpenStudy (anonymous):

ok that didnt work out ill just type it out

OpenStudy (anonymous):

with brackets then

OpenStudy (anonymous):

C=4d^-1/3b C=4(8500)^1/3(8.17) C=4(.049)(8.17) My question is how do they get (.049) from 8500^-1/3

OpenStudy (anonymous):

I have different numbers of course but Im just confused as to how they got the .049 if I can figure that part out I can solve my problem

OpenStudy (anonymous):

because 8500^(-1/3) simply equal 0.049 lol, if you put them on calculator, make sure you have brackets < http://www.wolframalpha.com/input/?i=8500%5E%28-1%2F3%29 >

OpenStudy (anonymous):

ok thank you sorry

OpenStudy (anonymous):

no problems